OpenStudy (anonymous):

What is the best first step to solve the system using elimination? 6x – 5y = 4 9x + y = –17

OpenStudy (aravindg):

add them

OpenStudy (aravindg):

first step multiply 2nd equation with 5 then add

OpenStudy (aravindg):

u can solve for x

OpenStudy (anonymous):

Yeah. That's what I'd do. You want the coefficient for y to cancel. Or you could choose to cancel the coeffecient for x, but in this case, that is a bit trickier.

hero (hero):

Solve both equations for y, set y = y, then solve for x, then solve for y.

OpenStudy (anonymous):

Practice makes a man perfact

OpenStudy (anonymous):

6x – 5y -4 = 9x + y +17

OpenStudy (anonymous):

6x-5y=4 9x+y=-17 multiply by 5 => 6x -5y=4 45x +5y=-85 (now u add them) ------------ 51x +0 =-81 => x=-81/51=-27/17 6*-27/17-5y=4=> y=46/17